OpenStudy (anonymous):

need EMERGENCY help for AP Calc BC exam tomorrow.. can't remember how to take derivative of inverse trig functions. (d/dx[sin^-1(x)]) HELP PLEASE!

OpenStudy (anonymous):

at this point that's what I'm thinking.

OpenStudy (anonymous):

:/ wish i could help

OpenStudy (anonymous):

no worries, thanks!

OpenStudy (across):

がんばって

OpenStudy (anonymous):

thanks, I'm gonna need it.

OpenStudy (anonymous):

d/dx[sin^-1(x)] = 1\sqrt[1 - x^2] Here's the proof: Let sin^-1(x) = y, therefore x = sin (y). Differentiate both sides with respect to x and we get 1 = d/dy(sin (y)) dy/dx --> 1 = cos (y) dy/dx --> dy/dx = 1/cos (y) Recall the elementary trig identity: sin^2 (y) + cos^2(y) = 1. Solve for cos (y) and we have: cos (y) = sqrt[1 - sin^2(y)] Recall sin (y) = x therefore cos (y) = sqrt[1 - x^2] and: d/dx arcsin (x) = 1 / sqrt[1 - x^2] Hope this helps!! :D

OpenStudy (anonymous):

THANK YOU SO MUCH! you're a lifesaver (: